# Medical Interpretability and Knowledge Maps of Large Language Models

### v1

@@ Abstract
We present a systematic study of medical-domain interpretability in Large Language Models (LLMs). We study how the LLMs both represent and process medical knowledge through four different interpretability techniques: (1) UMAP projections of intermediate activations, (2) gradient-based saliency with respect to the model weights, (3) layer lesioning/removal and (4) activation patching. We present knowledge maps of five LLMs which show, at a coarse-resolution, where knowledge about patient’s ages, medical symptoms, diseases and drugs is stored in the models. In particular for Llama3.3-70B, we find that most medical knowledge is processed in the first half of the model’s layers. In addition, we find several interesting phenomena: (i) age is often encoded in a non-linear and sometimes discontinuous manner at intermediate layers in the models, (ii) the disease progression representation is non-monotonic and circular at certain layers of the model, (iii) in Llama3.3-70B, drugs cluster better by medical specialty rather than mechanism of action, especially for Llama3.3-70B and (iv) Gemma3-27B and MedGemma-27B have activations that collapse at intermediate layers but recover by the final layers. These results can guide future research on fine-tuning, un-learning or de-biasing LLMs for medical tasks by suggesting at which layers in the model these techniques should be applied. We attached our source code to the paper for reproducibility.
@@ 2 Methods
UMAP: We use the Uniform Manifold Approximation and Projection (UMAP) McInnes et al. (2018b) to project the intermediate activations of the LLM into a 2D space for visualization. To get a quantitative measure of clustering, we compute the Silhuette Shahapure & Nicholas (2020) (cluster separation) score for each layer, which is a measure of how well the activations cluster in the lower-dimensional space. The Silhuette score is defined as:
@@ 3 Results
Figure 2: LLM Map for Llama3.3-70B showing where medical knowledge about age, symptoms, diseases, drugs and drug dosages is stored in the model. Each interval shown is estimated quantitatively from four types of analyses: clutering structure in UMAP embeddings, high weight saliency values, high degradation upon layer lesioning/ablation and high patching effect in activation patching.
@@
Disease Progression CSFS $\downarrow$ 3.25 ± 0.43 11 4.00 ± 1.22 16 4.00 ± 1.22 15 3.25 ± 0.43 13 4.00 ± 1.73 5
Disease Progression CSLS $\uparrow$ 7.00 ± 0.00 78 6.50 ± 0.87 48 6.25 ± 0.83 37 7.00 ± 0.00 62 7.00 ± 0.00 34
@@
Disease Progression CSFS $\downarrow$ 6.17 ± 1.95 6.10 ± 2.10 5.95 ± 2.11 5.73 ± 1.78 6.22 ± 1.94
Disease Progression CSLS $\uparrow$ 6.11 ± 1.42 4.79 ± 1.91 4.69 ± 1.89 6.43 ± 1.13 5.96 ± 1.46
@@ Appendix B Additional performance metrics
Disease Progression CSFS $\downarrow$ 7.25 ± 1.09 5.50 ± 2.60 5.50 ± 2.60 6.50 ± 0.87 5.50 ± 2.06
Disease Progression CSLS $\uparrow$ 6.50 ± 0.50 4.75 ± 2.28 4.75 ± 2.28 5.75 ± 1.09 5.50 ± 2.06

### v2

@@ Abstract
We present a systematic study of medical-domain interpretability in Large Language Models (LLMs). We study how the LLMs both represent and process medical knowledge through four different interpretability techniques: (1) UMAP projections of intermediate activations, (2) gradient-based saliency with respect to the model weights, (3) layer lesioning/removal and (4) activation patching. We present knowledge maps of five LLMs which show, at a coarse-resolution, where knowledge about patient’s ages, medical symptoms, diseases and drugs is stored in the models. In particular for Llama3.3-70B, we find that most medical knowledge is processed in the first half of the model’s layers. In addition, we find several interesting phenomena: (i) age is often encoded in a non-linear and sometimes discontinuous manner at intermediate layers in the models, (ii) the disease progression representation is non-monotonic and circular at certain layers of the model, (iii) in Llama3.3-70B, drugs cluster better by medical specialty rather than mechanism of action, especially for Llama3.3-70B and (iv) Gemma3-27B and MedGemma-27B have activations that collapse at intermediate layers but recover by the final layers. These results can guide future research on fine-tuning, un-learning or de-biasing LLMs for medical tasks by suggesting at which layers in the model these techniques should be applied. Source code is available at https://github.com/TheLumos/medical-interpretability-llms.
@@ 2 Methods
UMAP: We use the Uniform Manifold Approximation and Projection (UMAP) McInnes et al. (2018b) to project the intermediate activations of the LLM into a 2D space for visualization. We use n_neighbors=15n\_neighbors=15, min_dist=0.1min\_dist=0.1 and a cosine distance metric to build the UMAP connectivity graph. To get a quantitative measure of clustering, we compute the Silhuette Shahapure & Nicholas (2020) (cluster separation) score for each layer, which is a measure of how well the activations cluster in the lower-dimensional space. The Silhuette score is defined as:
@@ 3 Results
Figure 2: LLM Map for Llama3.3-70B showing where medical knowledge about age, symptoms, diseases, drugs and drug dosages is stored. Each interval shown is estimated quantitatively from four analyses: (1) clutering structure in UMAP embeddings, (2) high weight saliency values, (3) high degradation upon layer lesioning/ablation and (4) high patching effect in activation patching.
@@
Disease Progression CSFS $\downarrow$ 6.25 ± 2.17 0 4.25 ± 1.09 0 4.25 ± 1.09 0 4.25 ± 1.09 2 5.00 ± 1.58 3
Disease Progression CSLS $\uparrow$ 5.50 ± 2.06 74 5.75 ± 1.30 54 5.75 ± 1.30 52 6.75 ± 0.43 44 6.75 ± 0.43 29
@@
Disease Progression CSFS $\downarrow$ 6.96 ± 1.77 5.90 ± 2.14 5.72 ± 2.13 7.21 ± 1.59 6.58 ± 1.88
Disease Progression CSLS $\uparrow$ 5.08 ± 2.00 4.68 ± 2.10 4.66 ± 2.07 5.23 ± 1.98 5.65 ± 1.47
@@
We also computed quantitative metrics on six other medical and life science LLMs: OpenBioLLM-70B, PMC-LLaMA-13B, ClinicalCamel-70B, PalmyraMed-70B, Meditron-70B, HuatuoGPT-70B. We show the metrics for the best layer in Table 3, averages across all layers in Supplementary Table 7 and the last layer metrics in Supplementary Table 8.
Table 3: Per-concept interpretability metrics across medical LLMs. Values are best layer scores (max across depth) with the layer index.
Concept Metric Llama3 OpenBioLLM 70B PMC LLaMA 13B Clinical Camel 70B Palmyra Med 70B Meditron 70B HuatuoGPT o1 70B
Score Layer Score Layer Score Layer Score Layer Score Layer Score Layer
Age R2 (linear) $\uparrow$ 1.00 ± 0.00 16 0.99 ± 0.00 5 0.99 ± 0.00 38 1.00 ± 0.00 3 0.99 ± 0.00 20 1.00 ± 0.00 3
Symptoms Silhouette $\uparrow$ 0.22 ± 0.04 51 0.16 ± 0.04 30 0.15 ± 0.05 58 0.20 ± 0.04 47 0.16 ± 0.04 68 0.18 ± 0.04 50
Diseases Silhouette $\uparrow$ 0.07 ± 0.04 61 0.12 ± 0.04 26 0.15 ± 0.04 10 0.16 ± 0.04 54 0.16 ± 0.05 52 0.11 ± 0.04 59
Disease Progression CSFS $\downarrow$ 5.50 ± 2.60 0 4.25 ± 0.83 9 4.00 ± 0.71 19 5.75 ± 2.38 20 5.25 ± 1.48 19 6.25 ± 2.17 0
Disease Progression CSLS $\uparrow$ 7.00 ± 0.00 34 6.50 ± 0.50 38 6.75 ± 0.43 64 5.50 ± 2.06 72 6.75 ± 0.43 71 5.50 ± 2.06 74
Drugs Silhouette (mech.) $\uparrow$ 0.05 ± 0.06 48 0.09 ± 0.05 28 0.14 ± 0.07 50 0.09 ± 0.05 24 0.13 ± 0.07 51 0.11 ± 0.05 41
Drugs Silhouette (spec.) $\uparrow$ 0.15 ± 0.10 62 0.13 ± 0.07 27 0.22 ± 0.10 47 0.23 ± 0.10 47 0.23 ± 0.11 45 0.26 ± 0.11 46
Dosages Patching Effect $\uparrow$ 3.22 ± 1.20 79 16.00 ± 31.30 0 21.25 ± 46.75 1 0.26 ± 0.56 79 1.78 ± 3.32 49 1.99 ± 1.03 9
@@ 4 Discussion
4.1 Actionable recommendations
Removing age discontinuities and enforcing linear manifold for ages: If one wants to de-bias the age in order to remove the discontinuity at 18 years or to make the age manifold entirely linear, one can fine-tune only those exact layers where age is processed and use an age regularizer in the loss function that maintains linearity, such as $\mathcal{L_{\text{age}}}=\lambda*(1-\text{R}(\text{ages},\text{UMAP}(\text{prompt}(\text{ages}))))$, where R is the linear regression coefficient which will be 1 if the manifold of age embeddings is entirely linear.
[…]
Enforcing monotonicity in disease progression: During training, AI labs can add a regularizer that enforces monotonic disease progression, such as:
$\mathcal{R_{\text{monotonic}}}=\lambda\sum_{s=1}^{S-1}\left[\max\left(0,\;d(h_{s},h_{1})-d(h_{s+1},h_{1})\right)\right]^{2}$
which enforces the hidden-state activations $h_{s}$ at disease stage $s$ to move further away from $h_{1}$, the activation for stage 1 (healthy), where $d(\cdot,\cdot)$ is a chosen distance function. These regularizerscan be made even more complex by anchoring not just at $h_{1}$, but at other intermediate stages as well.
@@ 4.1 Actionable recommendations
Acknowledgments
This study was supported by Lumos AI. Computational credits were provided by Google Cloud.
@@ Appendix B Additional performance metrics
Disease Progression CSFS $\downarrow$ 7.50 ± 1.12 5.00 ± 1.22 5.00 ± 1.22 6.25 ± 1.48 7.25 ± 2.05
Disease Progression CSLS $\uparrow$ 4.50 ± 1.66 5.00 ± 2.12 4.00 ± 1.87 4.50 ± 1.66 5.00 ± 1.87
@@ Appendix D Full plots for all key results in main paper
Appendix G Results on other medical and life science LLMs
Concept Metric Llama3 OpenBioLLM 70B PMC LLaMA 13B Clinical Camel 70B Palmyra Med 70B Meditron 70B HuatuoGPT o1 70B
Age R2 (linear) $\uparrow$ 0.94 ± 0.07 0.87 ± 0.12 0.92 ± 0.10 0.97 ± 0.06 0.90 ± 0.12 0.98 ± 0.06
Symptoms Silhouette $\uparrow$ 0.09 ± 0.13 0.10 ± 0.06 0.06 ± 0.09 0.09 ± 0.09 0.07 ± 0.10 0.08 ± 0.08
Diseases Silhouette $\uparrow$ -0.01 ± 0.07 0.04 ± 0.05 0.06 ± 0.06 0.06 ± 0.07 0.06 ± 0.06 0.04 ± 0.06
Disease Progression CSFS $\downarrow$ 6.74 ± 1.91 6.99 ± 1.30 6.89 ± 1.30 6.82 ± 1.88 6.87 ± 1.24 7.11 ± 1.64
Disease Progression CSLS $\uparrow$ 5.26 ± 2.02 5.16 ± 1.85 5.61 ± 1.57 4.98 ± 1.98 5.66 ± 1.57 5.13 ± 2.01
Drugs Silhouette (mech.) $\uparrow$ -0.02 ± 0.06 0.03 ± 0.05 0.04 ± 0.07 0.04 ± 0.05 0.03 ± 0.07 0.05 ± 0.05
Drugs Silhouette (spec.) $\uparrow$ 0.03 ± 0.10 0.05 ± 0.07 0.07 ± 0.10 0.10 ± 0.10 0.07 ± 0.11 0.13 ± 0.11
Dosages Patching Effect $\uparrow$ 3.10 ± 1.15 -3.43 ± 12.65 12.06 ± 25.75 -0.16 ± 0.67 1.69 ± 2.81 1.70 ± 0.92
Concept Metric Llama3 OpenBioLLM 70B PMC LLaMA 13B Clinical Camel 70B Palmyra Med 70B Meditron 70B HuatuoGPT o1 70B
Age R2 (linear) $\uparrow$ 0.98 ± 0.01 0.91 ± 0.07 0.93 ± 0.05 0.98 ± 0.02 0.86 ± 0.12 0.98 ± 0.01
Symptoms Silhouette $\uparrow$ -0.07 ± 0.05 0.02 ± 0.04 -0.12 ± 0.03 -0.10 ± 0.04 -0.02 ± 0.04 -0.11 ± 0.04
Diseases Silhouette $\uparrow$ -0.07 ± 0.03 0.00 ± 0.04 -0.06 ± 0.04 -0.03 ± 0.04 -0.01 ± 0.04 -0.06 ± 0.04
Disease Progression CSFS $\downarrow$ 7.50 ± 1.12 7.25 ± 1.48 7.75 ± 2.17 7.00 ± 1.87 7.25 ± 1.30 7.50 ± 1.12
Disease Progression CSLS $\uparrow$ 5.25 ± 1.92 6.25 ± 0.43 6.50 ± 0.50 4.50 ± 1.66 6.50 ± 0.50 4.50 ± 1.66
Drugs Silhouette (mech.) $\uparrow$ 0.02 ± 0.06 0.05 ± 0.05 0.07 ± 0.07 0.07 ± 0.05 0.08 ± 0.07 0.09 ± 0.05
Drugs Silhouette (spec.) $\uparrow$ 0.07 ± 0.10 0.08 ± 0.07 0.09 ± 0.10 0.15 ± 0.10 0.11 ± 0.11 0.17 ± 0.11
Dosages Patching Effect $\uparrow$ 3.22 ± 1.20 2.75 ± 3.97 11.89 ± 24.77 0.26 ± 0.56 1.70 ± 2.74 1.78 ± 0.94